# Analysis of a class of Recursive Distributional Equations Including the Resistance of the Series-parallel graph

### v2

@@ B.1. Proof of Theorem 3
$\nu(N)=\sup\left\{G^{(N)}_{n}(y)-\varphi(\delta^{(N)}y,N^{-1}n)\,\mid\,y\in\mathbb{R},\,\,N^{-1}n\in[t_{0}-r,t_{0}+r]\right\}.$
Notice that (89), (84), and (83) imply that there is an $R>0$ such that, for any $N$,
$\sup\left\{G^{(N)}_{n}(y)-\varphi(\delta^{(N)}y,N^{-1}n)\,\mid\,\delta^{(N)}|y|\geq R,\,\,N^{-1}n\in[t_{0}-r,t_{0}+r]\right\}<0.$
Therefore, by the equation (86) relating $\{G^{(N)}_{n}\}$ to $G^{\star}$, it follows that there is a sequence $\{(N_{k},y_{k},n_{k})\}$ such that $(N^{-1}_{k},\nu(N_{k}))\to(0,0)$ as $k\to\infty$; $\delta^{(N)}|y_{k}|\leq R$ and $|N^{-1}n_{k}-t_{0}|\leq r$ for each $k$; and
(90) $G^{(N_{k})}_{n_{k}}(y_{k})-\varphi(\delta^{(N_{k})}y_{k},N_{k}^{-1}n_{k})=\nu(N_{k})\quad\text{for each}\,\,k.$
Further, since the set $\{(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r]\,\mid\,G^{\star}(x,t)=\varphi(x,t)\}$ is contained in $(-\infty,x_{0}]\times\{t_{0}\}$, there is no loss of generality in assuming that there is an $x^{\prime}_{0}\leq x_{0}$ such that $(\delta^{(N_{k})}y_{k},N_{k}^{-1}n_{k})\to(x^{\prime}_{0},t_{0})$ and $G^{\star}(x^{\prime}_{0},t_{0})=\varphi(x^{\prime}_{0},t_{0})$. Note, in particular, that the inequality $x^{\prime}_{0}\leq x_{0}$ implies, by the constancy of $\varphi_{1}$ in $(-\infty,x_{0}]$,
(91) $\lim_{k\to\infty}(\varphi_{1}(\delta^{(N_{k})}y_{k}),\partial_{x}\varphi_{1}(\delta^{(N_{k})}y_{k}),\partial_{x}^{2}\varphi_{1}(\delta^{(N_{k})}y_{k}))=(\varphi_{1}(x_{0}),0,0).$
Finally, for notational ease, let $\Phi_{k}:\mathbb{R}\times[t_{0}-r,t_{0}+r]\to\mathbb{R}$ be the function $\Phi_{k}(y,t)=\varphi(\delta^{(N_{k})}y,t)+\nu(N_{k})$. Since $\nu(N_{k})\to 0$, the bounds (89) imply that $0\leq\Phi_{k}\leq 1$ pointwise for all $k$ large enough. That is, $\Phi_{k}\in CDF(\overline{\mathbb{R}})$ for large enough $k$. By the choice of $\nu(N_{k})$,
$G^{(N_{k})}_{n}\leq\Phi_{k}(\cdot,N_{k}^{-1}n)\quad\text{pointwise for each}\,\,n\in[N(t_{0}-r),N(t_{0}+r)].$
@@
$T^{(N_{k})}G^{(N_{k})}_{n_{k}-1}\leq T^{(N_{k})}\Phi_{k}(\cdot,N_{k}^{-1}(n_{k}-1)).$
By (85), (90), and the definition of $\Phi_{k}$, this implies
$\Phi_{k}(y_{k},N_{k}^{-1}n_{k})=G^{(N_{k})}_{n_{k}}(y_{k})\leq T^{(N_{k})}G^{(N_{k})}_{n_{k}-1}(y_{k})\leq T^{(N_{k})}\Phi_{k}(y_{k},N_{k}^{-1}n_{k}-N_{k}^{-1}).$
Upon rearranging, this becomes
$N_{k}\Big(\Phi_{k}(y_{k},N_{k}^{-1}n_{k})-\Phi_{k}(y_{k},N_{k}^{-1}n_{k}-N_{k}^{-1})\Big)$
$\quad\quad\quad\leq N_{k}(\mathscr{L}^{(N_{k})}\Phi_{k})(y_{k},N_{k}^{-1}(n_{k}-1))+N_{k}Q^{(N_{k})}(\Phi_{k}(y_{k},N_{k}^{-1}(n_{k}-1)).$
By assumption (iii), the definition of $\Phi_{k}$, and the form of $\varphi$,
$(\mathscr{L}^{(N_{k})}\Phi_{k})(y_{k},N_{k}^{-1}(n_{k}-1))=(\mathscr{L}^{(N_{k})}(\varphi_{1})_{\delta^{(N_{k})}})(y_{k}),$
hence, by assumptions (v) and (vi), (91), and the form of $\varphi$, sending $k\to\infty$ above yields
$\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(0,0)+Q(\varphi(x_{0},t_{0})))$ $=\partial_{t}\varphi_{2}(t_{0})-\lim_{k\to\infty}\mathcal{F}(\partial_{x}\varphi_{1}(\delta^{(N_{k})}y_{k}),\partial_{x}^{2}\varphi_{1}(\delta^{(N_{k})}y_{k}))$
$\quad+\lim_{k\to\infty}N_{k}Q^{(N_{k})}(\varphi(\delta^{(N_{k})}y_{k},N_{k}^{-1}n_{k}))$
@@
Since $\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0$ by assumption, this yields (87) as desired.
@@ B.2. Proof of Corollary 4
In the proof that follows, it it convenient to work with the half-relaxed limits $\liminf_{*}F_{N}$ and $\limsup_{*}F_{N}$ of the rescaled CDF $F_{N}$ defined by
@@
$(\limsup\nolimits_{*}F_{N})(x,t)$ $=\lim_{\delta\downarrow 0}\sup\left\{F_{\epsilon}(x^{\prime},s)\,\mid\,|x^{\prime}-x|+|s-t|+N^{-1}\leq\delta\right\}.$
@@
$\underline{F}\leq\liminf\nolimits_{*}F_{\epsilon}\leq\limsup\nolimits^{*}F_{\epsilon}\leq\overline{F}\quad\text{in}\,\,\mathbb{R}\times[0,\infty).$
@@
where $\rho:(0,\infty)\to(0,\infty)$ is a smooth function supported in $[1,2]$ with $\int_{0}^{\infty}\rho(y)\,dy=1$. From the fact that $\rho$ is supported in $[1,2]$ and $F_{N}$ is nondecreasing in $x$, it follows that $F_{N,M}(x,0)\leq F_{N,M+1}(x,0)\leq F_{N}(x,0)$ for every $x\in\mathbb{R}$. Due to the monotonicity assumption (i), this implies that $F_{N,M}\leq F_{N,M+1}\leq F_{N}$ in $\mathbb{R}\times[0,\infty)$ for each $N$ and $M$. By Theorem 3, $F_{N,M}\to F_{M}$ locally uniformly in $\mathbb{R}\times[0,+\infty)$ as $N\to+\infty$, where $F_{M}$ is the solution of (61) with initial datum $F_{\text{in},M}$ given by $F_{\text{in},M}(x)=2^{M}\int_{0}^{\infty}F_{\text{in}}(x-y)\rho(2^{M}y)\,dy$. Further, by construction, the bound $F_{M}\leq F_{M+1}\leq\liminf\nolimits_{*}F_{N}$ holds pointwise in $\mathbb{R}\times[0,\infty)$ for any fixed $M$.
@@
satisfies $\partial_{t}F_{\infty}-\mathcal{F}(\partial_{x}F_{\infty},\partial_{x}^{2}F_{\infty})+Q(F_{\infty})\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and $F_{\infty}(\cdot,0)\geq F_{M}(\cdot,0)$ for any $M$. In particular, by the comparison principle, $F_{\infty}$ is at least as large as any viscosity subsolution of the initial-value problem (61). Therefore, $F_{\infty}\geq\underline{F}$ in $\mathbb{R}\times[0,\infty)$, and this implies $\liminf\nolimits_{*}F_{N}\geq F_{\infty}\geq\underline{F}$ as desired. ∎
@@ B.3. Lipschitz Estimate
Appendix C Effective Coefficients in the Series-Parallel Graph
This appendix computes the constants $\sigma_{D}$ and $a_{R}$ for the logarithms of the distance $D^{(n)}$ and the resistance $R^{(n)}$ of the critical series-parallel graph. In either case, the relevant forcing $f$ takes the form
@@ Appendix C Effective Coefficients in the Series-Parallel Graph
The left-inverse $g$ is given by
@@
Since $g(s)+f(g(s))=s$ for $s\geq\log 2$, the formulas from Section 1.4.1 become
$\sigma_{D}$ $=-\int_{0}^{\log 2}s\,ds-\int_{\log 2}^{\infty}\log\left(\frac{1}{1-e^{-s}}\right)\,ds,$
$a_{R}$ $=\int_{0}^{\log 2}s^{2}\,ds+\int_{\log 2}^{\infty}\log^{2}\left(\frac{1}{1-e^{-s}}\right)\,ds$
$\quad+2\int_{0}^{\log 2}s^{2}\,ds+2\int_{\log 2}^{\infty}s\log\left(\frac{1}{1-e^{-s}}\right)\,ds$
The integrals above can be computed using polylogarithms, a family of functions related to the Riemann zeta function $\zeta$. The dilogarithm and trilogarithm $Li_{2}$ and $Li_{3}$ are defined for $t\in[0,1]$ by the formulas
@@
To compute the integrals, use the change-of-variables $x=e^{-s}$, $ds=\frac{1}{x}dx$. In the case of $\sigma_{D}$, this leads to
$\sigma_{D}$ $=\int_{\frac{1}{2}}^{1}\log(x)\frac{1}{x}\,dx+\int_{0}^{\frac{1}{2}}\log(1-x)\frac{1}{x}\,dx$
@@
Finally, integrating-by-parts again and invoking the definitions of $Li_{2}$ and $Li_{3}$ yields
$a_{R}$ $=2\int_{0}^{1}Li_{2}(x)\frac{1}{x}\,dx-Li_{2}(1)\log(1)=2\int_{0}^{1}Li_{2}(x)\frac{1}{x}\,dx=2Li_{3}(1)=2\zeta(3).$

### v3

@@ B.1. Proof of Theorem 3
$\nu(N)=\sup\left\{G^{(N)}_{n}(y)-\psi(\delta^{(N)}y,N^{-1}n)\,\mid\,y\in\mathbb{R},\,\,N^{-1}n\in[t_{0}-r,t_{0}+r]\right\}.$
Notice that (89), (84), and (83) imply that there is an $R>0$ such that, for any large enough $N$,
$\sup\left\{G^{(N)}_{n}(y)-\psi(\delta^{(N)}y,N^{-1}n)\,\mid\,\delta^{(N)}|y|\geq R,\,\,N^{-1}n\in[t_{0}-r,t_{0}+r]\right\}<0.$
Therefore, by the equation (86) relating $\{G^{(N)}_{n}\}$ to $G^{\star}$, it follows that there is a sequence $\{(N_{k},y_{k},n_{k})\}$ such that $(N^{-1}_{k},\nu(N_{k}))\to(0,0)$ as $k\to\infty$; $\delta^{(N_{k})}|y_{k}|\leq R$ and $|N_{k}^{-1}n_{k}-t_{0}|\leq r$ for each $k$; and
(90) $G^{(N_{k})}_{n_{k}}(y_{k})-\psi(\delta^{(N_{k})}y_{k},N_{k}^{-1}n_{k})=\nu(N_{k})\quad\text{for each}\,\,k.$
Further, since the set $\{(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r]\,\mid\,G^{\star}(x,t)=\psi(x,t)\}$ is contained in $(-\infty,x_{0}]\times\{t_{0}\}$, there is no loss of generality in assuming that there is an $x^{\prime}_{0}\leq x_{0}$ such that $(\delta^{(N_{k})}y_{k},N_{k}^{-1}n_{k})\to(x^{\prime}_{0},t_{0})$ and $G^{\star}(x^{\prime}_{0},t_{0})=\psi(x^{\prime}_{0},t_{0})$. Note, in particular, that the inequality $x^{\prime}_{0}\leq x_{0}$ implies, by the constancy of $\psi_{1}$ in $(-\infty,x_{0}]$,
(91) $\lim_{k\to\infty}(\psi_{1}(\delta^{(N_{k})}y_{k}),\partial_{x}\psi_{1}(\delta^{(N_{k})}y_{k}),\partial_{x}^{2}\psi_{1}(\delta^{(N_{k})}y_{k}))=(\psi_{1}(x_{0}),0,0).$
Finally, for notational ease, let $\Psi_{k}:\mathbb{R}\times[t_{0}-r,t_{0}+r]\to\mathbb{R}$ be the function $\Psi_{k}(y,t)=\psi(\delta^{(N_{k})}y,t)+\nu(N_{k})$. Since $\nu(N_{k})\to 0$, the bounds (89) imply that $0\leq\Psi_{k}\leq 1$ pointwise for all $k$ large enough. That is, $\Psi_{k}\in CDF(\overline{\mathbb{R}})$ for large enough $k$. By the choice of $\nu(N_{k})$,
$G^{(N_{k})}_{n}\leq\Psi_{k}(\cdot,N_{k}^{-1}n)\quad\text{pointwise for each integer}\,\,n\in[N(t_{0}-r),N(t_{0}+r)].$
@@
$T^{(N_{k})}G^{(N_{k})}_{n_{k}-1}\leq T^{(N_{k})}\Psi_{k}(\cdot,N_{k}^{-1}(n_{k}-1)).$
By (85), (90), and the definition of $\Psi_{k}$, this implies
$\Psi_{k}(y_{k},N_{k}^{-1}n_{k})=G^{(N_{k})}_{n_{k}}(y_{k})\leq T^{(N_{k})}G^{(N_{k})}_{n_{k}-1}(y_{k})\leq T^{(N_{k})}\Psi_{k}(y_{k},N_{k}^{-1}n_{k}-N_{k}^{-1}).$
After rearranging terms, this becomes
$N_{k}\Big(\Psi_{k}(y_{k},N_{k}^{-1}n_{k})-\Psi_{k}(y_{k},N_{k}^{-1}n_{k}-N_{k}^{-1})\Big)$
$\quad\quad\quad\leq N_{k}(\mathscr{L}^{(N_{k})}\Psi_{k})(y_{k},N_{k}^{-1}(n_{k}-1))-N_{k}Q^{(N_{k})}(\Psi_{k}(y_{k},N_{k}^{-1}(n_{k}-1)).$
By assumption (iii), the definition of $\Psi_{k}$, and the form of $\psi$,
$(\mathscr{L}^{(N_{k})}\Psi_{k})(y_{k},N_{k}^{-1}(n_{k}-1))=(\mathscr{L}^{(N_{k})}(\psi_{1})_{\delta^{(N_{k})}})(y_{k}),$
hence, by assumptions (v) and (vi) and (91), sending $k\to\infty$ above yields
$\partial_{t}\psi(x_{0},t_{0})-\mathcal{F}(0,0)+Q(\psi(x_{0},t_{0})))$ $=\partial_{t}\psi_{2}(t_{0})-\lim_{k\to\infty}\mathcal{F}(\partial_{x}\psi_{1}(\delta^{(N_{k})}y_{k}),\partial_{x}^{2}\psi_{1}(\delta^{(N_{k})}y_{k}))$
$\quad+\lim_{k\to\infty}N_{k}Q^{(N_{k})}(\psi(\delta^{(N_{k})}y_{k},N_{k}^{-1}n_{k}))$
@@
Since $\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0$ by assumption and $\partial^{m}_{t}\psi(x_{0},t_{0})=\partial_{t}^{m}\varphi(x_{0},t_{0})$ for $m\in\{0,1\}$, this yields (87) as desired.
@@ B.2. Proof of Corollary 4
In the proof that follows, it it convenient to work with the half-relaxed limits $\liminf_{*}F_{N}$ and $\limsup^{*}F_{N}$ of the rescaled CDF $F_{N}$ defined by
@@
$(\limsup\nolimits^{*}F_{N})(x,t)$ $=\lim_{\delta\downarrow 0}\sup\left\{F_{\epsilon}(x^{\prime},s)\,\mid\,|x^{\prime}-x|+|s-t|+N^{-1}\leq\delta\right\}.$
@@
$\underline{F}\leq\liminf\nolimits_{*}F_{N}\leq\limsup\nolimits^{*}F_{N}\leq\overline{F}\quad\text{in}\,\,\mathbb{R}\times[0,+\infty).$
@@
where $\rho:(0,+\infty)\to(0,+\infty)$ is a smooth function supported in $[1,2]$ with $\int_{0}^{\infty}\rho(y)\,dy=1$. From the fact that $\rho$ is supported in $[1,2]$ and $F_{N}$ is nondecreasing in $x$, it follows that $F_{N,M}(x,0)\leq F_{N,M+1}(x,0)\leq F_{N}(x,0)$ for every $x\in\mathbb{R}$. Due to the monotonicity assumption (i), this implies that $F_{N,M}\leq F_{N,M+1}\leq F_{N}$ in $\mathbb{R}\times[0,\infty)$ for each $N$ and $M$. By Theorem 3, $F_{N,M}\to F_{M}$ locally uniformly in $\mathbb{R}\times[0,+\infty)$ as $N\to+\infty$, where $F_{M}$ is the solution of (61) with initial datum $F_{\text{in},M}$ given by $F_{\text{in},M}(x)=2^{M}\int_{0}^{\infty}F_{\text{in}}(x-y)\rho(2^{M}y)\,dy$. Further, by construction, the bound $F_{M}\leq F_{M+1}\leq\liminf\nolimits_{*}F_{N}$ holds pointwise in $\mathbb{R}\times[0,+\infty)$ for any fixed $M$.
@@
satisfies $\partial_{t}F_{\infty}-\mathcal{F}(\partial_{x}F_{\infty},\partial_{x}^{2}F_{\infty})+Q(F_{\infty})\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$ and $F_{\infty}(\cdot,0)\geq F_{M}(\cdot,0)$ for any $M$. In the limit $M\to+\infty$, this last bound becomes
$F_{\infty}(x,0)\geq\lim_{\delta\downarrow 0}F_{\text{in}}(x-\delta).$
Therefore, by the comparison principle (Theorem 4), $F_{\infty}$ is at least as large as any viscosity subsolution of the initial-value problem (61). Therefore, $F_{\infty}\geq\underline{F}$ in $\mathbb{R}\times[0,+\infty)$, and this implies $\liminf\nolimits_{*}F_{N}\geq F_{\infty}\geq\underline{F}$ as desired. ∎
@@ B.3. Lipschitz Estimate
Appendix C Effective Coefficients in the Series-Parallel Graph Case
This appendix computes the constants $\sigma_{\mathcal{D}}$ and $a_{R}$ in the case of the logarithm of the distance and the resistance of the series-parallel graph. Recall that, in this setting, the relevant forcing $f$ takes the form
@@ Appendix C Effective Coefficients in the Series-Parallel Graph Case
The corresponding function $g$ from Section 1.4.1 is given by
@@
Since $g(s)+f(g(s))=s$ for $s\geq\log 2$, the formulae for $\sigma_{\mathcal{D}}$ and $a_{R}$ become
$\sigma_{\mathcal{D}}$ $=-\int_{0}^{\log 2}s\,ds-\int_{\log 2}^{\infty}\log\left(\frac{1}{1-e^{-s}}\right)\,ds,$
$a_{R}$ $=3\int_{0}^{\log 2}s^{2}\,ds+\int_{\log 2}^{\infty}\log^{2}\left(\frac{1}{1-e^{-s}}\right)\,ds+2\int_{\log 2}^{\infty}s\log\left(\frac{1}{1-e^{-s}}\right)\,ds$
The integrals above can be computed using polylogarithms, a family of functions related to the Riemann zeta function $\zeta$. The dilogarithm and trilogarithm $Li_{2}$ and $Li_{3}$ are defined for $t\in[0,1]$ via the integrals
@@
To compute the integrals, use the change-of-variables $x=e^{-s}$, $ds=\frac{1}{x}dx$. In the case of $\sigma_{\mathcal{D}}$, this leads to
$\sigma_{\mathcal{D}}$ $=\int_{\frac{1}{2}}^{1}\log(x)\frac{1}{x}\,dx+\int_{0}^{\frac{1}{2}}\log(1-x)\frac{1}{x}\,dx$
@@
Finally, integrating by parts and invoking the definitions of $Li_{2}$ and $Li_{3}$ yields
$a_{R}$ $=2\int_{0}^{1}Li_{2}(x)\frac{1}{x}\,dx-2Li_{2}(1)\log(1)=2\int_{0}^{1}Li_{2}(x)\frac{1}{x}\,dx=2Li_{3}(1)=2\zeta(3).$